Flying against the wind, an airplane travels 6300km in 7 hours. Flying with the wind, the same plane travels 11,200km in 8 hours. What is the rate of the plane in still air and what is the rate of the wind?

Answer by Boreal(15235)

The speed against the wind is 6300/7=900 km/h
With it is 11200/8=1400 km/h
x+w=1400
x-w=900
2x=2300
x=1150 km/h speed of plane
w=250 km/h speed of wind